# Blowin’ in the non-isothermal wind: core-powered mass loss with hydrodynamic radiative transfer

### v1

@@ 1 Introduction
Importantly, the outer atmospheres of sub-Neptunes need not be isothermal. For example, if the opacity to the infrared outgoing radiation, $\kappa_{\mathrm{P,therm}}$ is not the same as the opacity to the incident visible radiation, $\kappa_{\mathrm{P},\odot}$, then the upper atmosphere will have a more complex, non-isothermal temperature structure. This was first theorized to have important implications for the flared, upper atmospheres of protoplanetary disks due to favorable opacity functions of primitive dust (Calvet et al., 1991; Chiang & Goldreich, 1997; Dullemond et al., 2001) and can also occur in the upper atmospheres of planets, even when presumed to be dust-free (Hubeny et al., 2003; Guillot, 2010). This difference in opacities between spectral bands is often quantified as a ratio
@@
These non-isothermal temperature structures in the optically thin regions of atmospheres have been extensively studied in exoplanetary contexts (e.g. Burrows et al., 2007; Fortney et al., 2008; Ito et al., 2015; Mollière et al., 2015; Wyttenbach et al., 2015), due to their effects on observed transmission spectra and upper atmospheric chemical equilibrium. But by changing the density profile of the upper atmosphere, such non-isothermal profiles also affect mass loss rates (Schulik & Booth, 2023; Schulik et al., 2024). Variations in $\gamma$ are expected across the diverse temperatures, pressures, and compositions of the sub-Neptune population. In this work, we investigate whether these different potential values of $\gamma$ can affect the expected mass loss rates, and therefore the evolution, of planets undergoing core-powered mass loss.
@@
This paper is structured as follows. In Section 2, we present the atmospheric structure and evolution theory that form the basis of this work and describe our simulation methods. In Section 3, we present and analyze our atmospheric mass loss rate and evolution simulation results. In Section 4, we discuss our key assumptions and future possible directions, followed by a conclusion in Section 5.
@@ 2.1 Atmospheric structure theory
The basic picture of sub-Neptune composition and structure that we adopt in this work closely follows that of Misener & Schlichting (2021), which is one of a silicate-rich core surrounded by a hydrogen-rich envelope, motivated by extensive previous work (e.g. Lopez & Fortney, 2014; Ginzburg et al., 2016; Owen & Wu, 2017). This composition is supported by comparison of mass loss models to the radius valley’s location and its dependence on orbital period and stellar mass (Gupta & Schlichting, 2019; Rogers & Owen, 2021). In our model, an incompressible silicate core of mass $M_{\mathrm{p}}$ extends to a core radius $R_{\mathrm{c}}/R_{\oplus}\simeq(M_{\mathrm{p}}/M_{\oplus})^{1/4}$ (Valencia et al., 2006). Above the core, the hydrogen-rich envelope can be modeled to first order as a convective region at the base topped by a radiative region. The transition between these two regions occurs at the radiative-convective boundary (RCB) radius, $R_{\mathrm{rcb}}$. In this work, we ignore the effects of compositional mixing between silicates and hydrogen that can lead to non-convective regions deep within sub-Neptunes (Misener & Schlichting, 2022; Markham et al., 2022).
@@
where $\rho_{\mathrm{rcb}}=\rho(r=R_{\mathrm{rcb}})$ is the density at the radiative-convective boundary and $\gamma_{\mathrm{ad}}$ is the adiabatic index of the atmosphere (not to be confused with the ratio of the visible and infrared opacities, which will be referred to as $\gamma$ throughout this work). We take $\gamma_{\mathrm{ad}}=7/5$, appropriate for diatomic hydrogen. The so-called ‘modified Bondi radius’ is defined for convenience as $R_{\mathrm{B}}^{\prime}\equiv(\gamma_{\mathrm{ad}}-1)/\gamma_{\mathrm{ad}}\times GM_{\mathrm{p}}\mu/(k_{\mathrm{B}}T_{\mathrm{rcb}})$, where $G$ is the gravitational constant, $\mu=2m_{\mathrm{p}}$ is the molecular weight of H2, $k_{\mathrm{B}}$ is the Boltzmann constant, and $T_{\mathrm{rcb}}=T(r=R_{\mathrm{rcb}})$ is the temperature at the radiative-convective boundary. Since the mass in the radiative region is typically negligible, the mass contained in the atmosphere, $M_{\mathrm{atm}}$, can be found by integrating Equation 2 over the convective region. Meanwhile, to calculate the planet’s total energy, we follow Misener & Schlichting (2021). The total energy is the sum of the available energy in the core and atmosphere, $E=E_{\mathrm{core}}+E_{\mathrm{atm}}$. Since we take the core to be incompressible, the core’s available energy is purely thermal: $E_{\mathrm{core}}=(1/(\gamma_{\mathrm{c}}-1))M_{\mathrm{p}}/\mu_{\mathrm{c}}k_{\mathrm{B}}T_{\mathrm{c}}$, where $\gamma_{\mathrm{c}}\sim 4/3$ is the core’s ‘adiabatic index’, i.e. a representation of its specific heat capacity (Scipioni et al., 2017), and $\mu_{\mathrm{c}}=60$ amu is the core’s mean molecular weight. We assume the core is isothermal and thermally coupled to the base of the atmosphere, such that $T_{\mathrm{c}}=T(r=R_{\mathrm{c}})$. The atmosphere’s energy is the sum of its gravitational potential and thermal energies:
@@
Equation 5 applies until radiative diffusion no longer dominates energy transport, i.e. the optical depth by the Rosseland mean opacity $\tau_{\mathrm{R}}<1$. In the results presented in this work, we use a double-gray radiative-transfer hydrodynamic model to self-consistently calculate the temperature and density profiles above this point, but it is useful to understand the key physical processes. Previous work has shown that to first order, the radiative atmosphere can be characterized by two temperature regimes (e.g. Guillot, 2010). In the bottom of this region, where the optical depth to outgoing blackbody radiation, $\tau_{\mathrm{P,therm}}>1$, the expected temperature $T_{\mathrm{in}}\approx T_{\mathrm{eq}}/2^{1/4}$. This inner region temperature (not to be confused with the ‘internal temperature’ representing the net outgoing heat flux, $T_{\mathrm{int}}$) is achieved in the limit that $\gamma\gg 1$ or $\gamma\ll 1$, while $T_{\mathrm{in}}$ is expected to be closer to $T_{\mathrm{eq}}$ when $\gamma\sim 1$ (Parmentier & Guillot, 2014). Meanwhile, we term the expected temperature in the outer atmosphere, where the optical depth to incident stellar radiation $\tau_{\mathrm{P},\odot}<1$, $T_{\mathrm{out}}$. This outer or skin temperature depends on the thermal and stellar opacities and their ratio and can be expressed as $T_{\mathrm{out}}=T_{\mathrm{eq}}(\gamma/4)^{1/4}$ (Guillot, 2010; Schulik & Booth, 2023). If the opacity to incoming stellar radiation is lower than the opacity to outgoing thermal radiation ($\gamma<1$), then the radiative upper atmosphere is better at cooling than at heating, and the temperature will be colder than the deeper regions ($T_{\mathrm{out}}<T_{\mathrm{in}}$). Conversely, if the opacity to incoming stellar radiation is higher than the opacity to outgoing thermal radiation ($\gamma>1$), then the upper atmosphere will absorb incident stellar light without being able to cool efficiently. Thus the outer atmosphere will be hotter than the deeper regions, causing a temperature inversion ($T_{\mathrm{out}}>T_{\mathrm{in}}$).
where $s=(r^{2}-R_{\mathrm{trans}}^{2})^{1/2}$ is the chord length.
In Figure 1, we summarize the key atmospheric physics described in this and the following sections. In each panel of the schematic, which is not to scale, we show example temperature profiles as functions of radii for three prototypical cases: an atmosphere with $\gamma\gg 1$, $\gamma\sim 1$, and $\gamma\ll 1$ from top to bottom. We show three key radii with dashed black vertical lines: the radiative-convective boundary radius $R_{\mathrm{rcb}}$, the transit radius $R_{\mathrm{trans}}$, and the sonic radius of the outflow $R_{\mathrm{s}}$ (see Eq. 11 below). We also show with dash-dotted lines the radii at which the optical depth reaches unity for the three opacities that determine the structure of the atmosphere: the thermal Rosseland optical depth $\tau_{\mathrm{R}}$ in blue, the thermal Planck optical depth $\tau_{\mathrm{P,therm}}$ in red, and the two-temperature stellar Planck optical depth $\tau_{\mathrm{P},\odot}$ in yellow. The domain of our hydrodynamic simulation is shown by the tan shading, extending outwards from its inner boundary $R_{\mathrm{domain}}$ (see Sec. 2.6 below for details). Finally, we show the key temperatures defined in this section with gray horizontal dotted lines: the equilibrium temperature $T_{\mathrm{eq}}$, temperature at the radiative-convective boundary $T_{\mathrm{rcb}}$, temperature in the inner radiative region $T_{\mathrm{in}}$, and temperature in the outer radiative region $T_{\mathrm{out}}$.
@@ 2.2 Opacities
The opacities are a combination of broadband H absorption and line absorption by molecules in the infrared and alkali metals and oxides in the visible. Therefore, depending on a variety of atmospheric processes, including temperature and density, as well as composition, the value of the ratio could vary significantly from planet to planet. We therefore test a range of values of the opacity to solar radiation $\kappa_{\mathrm{P},\odot}$ ranging from 0.225 to 225 cm2 g-1, corresponding to opacity ratios between 0.03 and 30, which encompass the range of temperature, density, and compositions relevant to sub-Neptunes (Freedman et al., 2014). In this way we remain agnostic to the actual opacity sources that cause the atmospheric effects we study here, and rather seek to quantify the magnitude of differences which are possible. In reality, using wavelength-resolved opacities are optimal, because the exact line structure can have dramatic effects on the mass loss rate (Schulik et al., 2024). In the future, once these physical effects are better understood, real opacity models could be used to perform bespoke modeling of the evolution of individual planets.
@@ 2.3 Atmospheric evolution theory
We use aiolos, a 1D radiation-hydrodynamics code first presented in Schulik & Booth (2023). We relegate most technical details to that paper, but we briefly restate the most important numerical approaches here. aiolos is a multispecies hydrodynamic code, solving one instance of the Euler system (Toro, 2009) per species via the HLLC Riemann solver (Toro et al., 1994), using the Käppeli & Mishra (2014) well-balancing scheme to stabilize hydrostatic regions against gravity. We only use one fluid in this work. The temperatures of our single gas species are computed in a double-grey approximation, with one radiation band representing the incoming and one the outgoing radiation. The incoming absorbed radiation is solved via the trivial solution of the Lambert-Beer law (Commerçon et al., 2011; Kuiper & Klessen, 2013), as we neglect scattering. The outgoing, thermal radiation, whose mean photon field density is coupled to the gas temperature field, is treated numerically via flux-limited diffusion (Levermore & Pomraning, 1981) with the flux-limiter from (Kley, 1989), and the radiation matrix is solved similar to the methods presented in (Bitsch et al., 2013; Lega et al., 2014).
@@ 2.6 Optically thick extrapolation
Beginning from the bottom of the aiolos domain, we increase the pressure by a small increment, $\Delta P=P/100$, and increment the temperature and radius according to Equations 5 and 6 respectively. We recalculate the gradients, and continue this until the Schwarzschild criterion is met, i.e. $\partial\ln T/\partial\ln P=2/7$. At this radius, the radiative-convective boundary, $R_{\mathrm{rcb}}$, we switch to an adiabatic gradient, calculated using the methods of Misener & Schlichting (2021). We integrate this adiabatic profile to infer an atmospheric mass $M_{\mathrm{atm}}$, often referred to as a fraction of the planet’s total mass, $f\equiv M_{\mathrm{atm}}/M_{\mathrm{p}}$. In this way, we connect an aiolos boundary condition to the atmospheric evolution state it represents.

### v2

@@ 1 Introduction
Importantly, the outer atmospheres of sub-Neptunes need not be isothermal. For example, if the opacity to the infrared outgoing radiation, $\kappa_{\mathrm{P,therm}}$ is not the same as the opacity to the incident visible radiation, $\kappa_{\mathrm{P},\odot}$, then the upper atmosphere will have a more complex, non-isothermal temperature structure, which can produce temperatures both hotter and colder than the equilibrium temperature in different regions. This was first theorized to have important implications for the flared, upper atmospheres of protoplanetary disks due to favorable opacity functions of primitive dust (Calvet et al., 1991; Chiang & Goldreich, 1997; Dullemond et al., 2001) and can also occur in the upper atmospheres of planets, even when presumed to be dust-free (Hubeny et al., 2003; Guillot, 2010). This difference in opacities between spectral bands is often quantified as a ratio
@@
These non-isothermal temperature structures in the optically thin regions of atmospheres have been extensively studied in exoplanetary contexts (e.g. Burrows et al., 2007; Fortney et al., 2008; Ito et al., 2015; Mollière et al., 2015; Wyttenbach et al., 2015), due to their effects on observed transmission spectra and upper atmospheric chemical equilibrium. But by changing the density profile of the upper atmosphere, such non-isothermal profiles also affect mass loss rates (Schulik & Booth, 2023; Schulik et al., 2024). Variations in $\gamma$ are expected across the diverse temperatures, pressures, and compositions of the sub-Neptune population. In this work, we investigate whether consideration of non-gray effects can affect the expected mass loss rates, and therefore the evolution, of planets undergoing core-powered mass loss. In doing so, we present the first test of core-powered mass loss using a hydrodynamic radiative transfer model.
@@
This paper is structured as follows. We begin with an overview of the key physics of upper atmospheric temperature structure and its expected effect on mass loss rates in Section 2. In Section 3, we present the atmospheric structure and evolution theory that form the basis of this work and describe our simulation methods. In Section 4, we present and analyze our atmospheric mass loss rate and evolution simulation results. In Section 5, we discuss our key assumptions and future possible directions, followed by a conclusion in Section 6.
2 Overview and Motivations
We first provide an overview of the key physical processes that determine, to first order, the upper atmospheric structures we analyze in this manuscript. Previous work has shown that using the double-gray approximation, the radiative atmospheres of planets can be characterized by two temperature regimes (e.g. Guillot, 2010). In the bottom of this region, where the optical depth to outgoing thermal radiation, $\tau_{\mathrm{P,therm}}>1$, the atmosphere reaches thermal equilibrium. Therefore the expected temperature in the inner region is close to the planetary equilibrium temperature: $T_{\mathrm{in}}\approx T_{\mathrm{eq}}/2^{1/4}$. This inner region temperature (not to be confused with the ‘internal temperature’ representing the net outgoing heat flux, $T_{\mathrm{int}}$) is achieved in the limit that $\gamma\gg 1$ or $\gamma\ll 1$, while $T_{\mathrm{in}}$ is expected to be closer to $T_{\mathrm{eq}}$ when $\gamma\sim 1$ (Parmentier & Guillot, 2014).
Meanwhile, we term the expected temperature in the outer atmosphere, where the optical depth to incident stellar radiation $\tau_{\mathrm{P},\odot}<1$, $T_{\mathrm{out}}$. In a fully gray atmosphere, the entire radiative region should be isothermal at the blackbody temperature, i.e. $T_{\mathrm{out}}\sim T_{\mathrm{in}}\sim T_{\mathrm{eq}}$. However, non-gray, i.e. wavelength-dependent, opacities can change the temperature profile. If the opacity to incoming stellar radiation is lower than the opacity to outgoing thermal radiation ($\gamma<1$), then the upper atmosphere radiates thermal energy away more efficiently than it absorbs incident stellar energy, and the temperature will be colder than the deeper regions ($T_{\mathrm{out}}<T_{\mathrm{in}}$). Conversely, if the opacity to incoming stellar radiation is higher than the opacity to outgoing thermal radiation ($\gamma>1$), then the upper atmosphere will absorb incident stellar light without being able to cool efficiently. Thus the outer atmosphere will be hotter than the deeper regions, causing a temperature inversion ($T_{\mathrm{out}}>T_{\mathrm{in}}$). Therefore, this outer, or skin, temperature scales with the ratio of the thermal and stellar opacities and can be expressed as $T_{\mathrm{out}}=T_{\mathrm{eq}}(\gamma/4)^{1/4}$ (Guillot, 2010; Schulik & Booth, 2023). In either case, due to steady state disequilibrium between the internal energy of the gas and the radiation field, an atmosphere with $\gamma\neq 1$ can have an outer temperature that deviates from the equilibrium temperature.
where $s=(r^{2}-R_{\mathrm{trans}}^{2})^{1/2}$ is the chord length.
In Figure 1, we summarize the key atmospheric physics described in this and the following sections. In each panel of the schematic, which is not to scale, we show example temperature profiles as functions of radii for three prototypical cases: an atmosphere with $\gamma\gg 1$, $\gamma\sim 1$, and $\gamma\ll 1$ from top to bottom. We show three key radii with dashed black vertical lines: the radiative-convective boundary (RCB) radius $R_{\mathrm{rcb}}$, the transit radius $R_{\mathrm{trans}}$, and the sonic radius of the outflow $R_{\mathrm{s}}$ (see Eq. 11 below). We also show with dash-dotted lines the radii at which the optical depth reaches unity for the three opacities that determine the structure of the atmosphere: the thermal Rosseland optical depth $\tau_{\mathrm{R}}$ in blue, the thermal Planck optical depth $\tau_{\mathrm{P,therm}}$ in red, and the two-temperature stellar Planck optical depth $\tau_{\mathrm{P},\odot}$ in yellow. The domain of our hydrodynamic simulation is shown by the tan shading, extending outwards from its inner boundary $R_{\mathrm{domain}}$ (see Sec. 3.6 below for details). Finally, we show the key temperatures defined in this section with gray horizontal dotted lines: the equilibrium temperature $T_{\mathrm{eq}}$, temperature at the radiative-convective boundary $T_{\mathrm{rcb}}$, temperature in the inner radiative region $T_{\mathrm{in}}$, and temperature in the outer radiative region $T_{\mathrm{out}}$.
These changes in temperature produced by opacity alterations can affect the mass loss rates. One important change is to the density profile in the outer atmosphere. To first order, the density profile follows hydrostatic equilibrium, falling off exponentially with the lapse rate proportional to the temperature. Therefore, in regions where the temperature is lower than the equilibrium temperature, $T<T_{\mathrm{eq}}$, the density decreases more quickly with increasing radius than in an isothermal atmosphere at $T_{\mathrm{eq}}$. Conversely, in regions with $T>T_{\mathrm{eq}}$, the density falls off more slowly than in the isothermal case. Increasing the temperature also increases the sound speed of the gas, leading to a more rapid acceleration out of the planet’s gravitational potential and a decrease in the sonic radius $R_{\mathrm{s}}$. To first order, these temperature effects lead to an increase in mass loss rate for planets with $\gamma\gg 1$ and a concomitant decrease in mass loss rate in $\gamma\ll 1$ cases. However, as we show in Section 4, the interaction between the various physical mechanisms described here can lead to additional complexity not captured by these simplified equations. We therefore employ a radiation-hydrodynamics code to model these atmospheres more self-consistently, which we describe in the following section.
@@ 3.1 Atmospheric structure theory
The basic picture of sub-Neptune composition and structure that we adopt in this work closely follows that of Misener & Schlichting (2021), which is one of a silicate-rich core surrounded by a hydrogen-rich envelope, motivated by extensive previous work (e.g. Lopez & Fortney, 2014; Ginzburg et al., 2016; Owen & Wu, 2017). This composition is supported by comparison of mass loss models to the radius valley’s location and its dependence on orbital period and stellar mass (Gupta & Schlichting, 2019; Rogers & Owen, 2021). In our model, an incompressible silicate core of mass $M_{\mathrm{p}}$ extends to a core radius $R_{\mathrm{c}}/R_{\oplus}\simeq(M_{\mathrm{p}}/M_{\oplus})^{1/4}$ (Valencia et al., 2006). Above the core, the hydrogen-rich envelope can be modeled to first order as a convective region at the base topped by a radiative region. The transition between these two regions occurs at the radiative-convective boundary radius, $R_{\mathrm{rcb}}$. In this work, we ignore the effects of compositional mixing between silicates and hydrogen that can lead to non-convective regions deep within sub-Neptunes (Misener & Schlichting, 2022; Markham et al., 2022) and to additional heat release due to condensation and demixing processes, including core formation and silicate rainout (e.g. Vazan et al., 2024).
@@
where $\rho_{\mathrm{rcb}}=\rho(r=R_{\mathrm{rcb}})$ is the density at the radiative-convective boundary and $\gamma_{\mathrm{ad}}$ is the adiabatic index of the atmosphere (not to be confused with the ratio of the visible and infrared opacities, which will be referred to as $\gamma$ throughout this work). We take $\gamma_{\mathrm{ad}}=7/5$, appropriate for diatomic hydrogen. The so-called ‘modified Bondi radius’ is defined for convenience as $R_{\mathrm{B}}^{\prime}\equiv(\gamma_{\mathrm{ad}}-1)/\gamma_{\mathrm{ad}}\times GM_{\mathrm{p}}\mu/(k_{\mathrm{B}}T_{\mathrm{rcb}})$, where $G$ is the gravitational constant, $\mu=2m_{\mathrm{p}}$ is the molecular weight of H2, $k_{\mathrm{B}}$ is the Boltzmann constant, and $T_{\mathrm{rcb}}=T(r=R_{\mathrm{rcb}})$ is the temperature at the radiative-convective boundary. Since the mass in the radiative region is typically negligible, the mass contained in the atmosphere, $M_{\mathrm{atm}}$, can be found by integrating Equation 3 over the convective region.
Meanwhile, to calculate the planet’s total energy, we follow Misener & Schlichting (2021). The total energy is the sum of the available energy in the core and atmosphere, $E=E_{\mathrm{core}}+E_{\mathrm{atm}}$. Since we take the core to be incompressible, the core’s available energy is purely thermal: $E_{\mathrm{core}}=(1/(\gamma_{\mathrm{core}}-1))M_{\mathrm{p}}/\mu_{\mathrm{core}}k_{\mathrm{B}}T_{\mathrm{core}}$, where $\gamma_{\mathrm{core}}\sim 4/3$ is the core’s ‘adiabatic index’, i.e. a representation of its specific heat capacity (Scipioni et al., 2017), and $\mu_{\mathrm{core}}=60$ amu is the core’s mean molecular weight. We assume the core is isothermal and thermally coupled to the base of the atmosphere, such that $T_{\mathrm{core}}=T(r=R_{\mathrm{c}})$. We neglect the energy released by differentiation in the interior, making our assumed core heat reservoir a conservative lower bound. However, the gravitational energy released by such processes will be at most the same order of magnitude as the formation energy of the planet, so our estimate is accurate to a factor of a few.
The atmosphere’s energy is the sum of its gravitational potential and thermal energies:
@@
Equation 6 applies until radiative diffusion no longer dominates energy transport, i.e. the optical depth by the Rosseland mean opacity $\tau_{\mathrm{R}}<1$. At this point, the temperature profiles depend strongly on the wavelength-dependent absorption and emission of light, and the physics described in Section 2 becomes paramount. To capture these complexities, in the results presented in this work we use a double-gray radiative-transfer hydrodynamic model to self-consistently calculate the temperature and density profiles above this point, which we describe in detail in Section 3.4 below.
@@ 3.2 Opacities
The opacities are a combination of broadband H absorption and line absorption by molecules in the infrared and alkali metals and oxides in the visible. Therefore, depending on a variety of atmospheric processes, including temperature and density, as well as composition, the value of the ratio could vary significantly from planet to planet. To understand the basic physics at play and make a first-order estimate of the magnitude of potential variation in mass loss rates, we adopt constant values of opacity for each run and iterate over a range of realistic opacities. Specifically, we test a range of values of the opacity to solar radiation $\kappa_{\mathrm{P},\odot}$ ranging from 0.225 to 225 cm2 g-1, corresponding to opacity ratios between 0.03 and 30, which encompass the range of temperature, density, and compositions relevant to sub-Neptunes (Freedman et al., 2014). In this way we remain agnostic to the actual opacity sources that cause the atmospheric effects we study here, and rather seek to quantify the magnitude of differences which are possible. In reality, using wavelength-resolved opacities is optimal, because the exact line structure can have dramatic effects on the mass loss rate (Schulik et al., 2024). In the future, once these physical effects are better understood, real opacity models could be used to perform bespoke modeling of the evolution of individual planets. We discuss potential variations in opacity both radially within the atmosphere and through time in Section 5.3.
@@ 3.3 Atmospheric evolution theory
We use aiolos, a 1D radiation-hydrodynamics code first presented in Schulik & Booth (2023). We relegate most technical details to that paper, but we briefly restate the most important numerical approaches here. aiolos is a multispecies hydrodynamic code, solving one instance of the Euler system (Toro, 2009) per species via the HLLC Riemann solver (Toro et al., 1994), using the Käppeli & Mishra (2014) well-balancing scheme to stabilize hydrostatic regions against gravity. We only use one fluid in this work. The temperatures of our single gas species are computed in a double-gray approximation, with one radiation band representing the incoming and one the outgoing radiation. The incoming absorbed radiation is solved via the trivial solution of the Lambert-Beer law (Commerçon et al., 2011; Kuiper & Klessen, 2013), as we neglect scattering. The outgoing, thermal radiation, whose mean photon field density is coupled to the gas temperature field, is treated numerically via flux-limited diffusion (Levermore & Pomraning, 1981) with the flux-limiter from (Kley, 1989), and the radiation matrix is solved similar to the methods presented in (Bitsch et al., 2013; Lega et al., 2014).
@@ 3.6 Optically thick extrapolation
Beginning from the bottom of the aiolos domain, we increase the pressure by a small increment, $\Delta P=P/100$, and increment the temperature and radius according to Equations 6 and 7 respectively. We recalculate the gradients, and continue this until the Schwarzschild criterion is met, i.e. $\partial\ln T/\partial\ln P=2/7$. At this radius, the radiative-convective boundary, $R_{\mathrm{rcb}}$, we switch to an adiabatic gradient, calculated using the methods of Misener & Schlichting (2021). We integrate this adiabatic profile to infer an atmospheric mass $M_{\mathrm{atm}}$, often referred to as a fraction of the planet’s total mass, $f\equiv M_{\mathrm{atm}}/M_{\mathrm{p}}$. In this way, we connect an aiolos boundary condition to the atmospheric evolution state it represents.